I need to find the foci, the center and the vertices of an ellipse
Equation: (x+50)^2/25 + (y-80)^2/10 < 1

Answer by lwsshak3(11628)

I need to find the foci, the center and the vertices of an ellipse
Equation:(x+50)^2/25 + (y-80)^2/10 = 1
given ellipse has a horizontal major axis
center: (-50,80)
a^2=25
a=5
vertices: (-50±a,80)=(-50±5,80)=(-55,80) and (-45,80)
b^2=10
c^2=a^2-b^2=25-10=15
c=√15≈3.87
foci: (-50±c,80)=(-50±3.87,80)=(-53.87,80) and (-46.13,80)